# Morse-Bott inequalities, Topology Change and Cobordisms to Nothing

### v1

@@ 4.2 Type IIB string theory on compact Riemann surfaces
    $\Delta V_{F}\approx-\left(\frac{|\hat{\tau}^{(0)}|^{2}}{\hat{\tau}_{2}^{(0)}}Q^{2}+\frac{1}{\hat{\tau}_{2}^{(0)}}Q^{\prime}{}^{2}\right)\,,$ (4.22)
@@
[…]
    $\mathbb{A}_{AB}H_{2}^{A}\wedge\star H_{2}^{B}$ $=\frac{\star 1}{\tau_{2}}|H_{2}^{c}-\tau H_{2}^{c^{\prime}}|^{2}+...$
    $=\frac{1}{\tau_{2}}|H_{2}^{-}|^{2}+\tau_{2}|H_{2}^{+}|^{2}\,,$ (4.23)
    where $\tau=\tau_{1}+i\tau_{2}$ is the factorized complex structure associated to $c$ and $c^{\prime}$ and the ()($\dots$)( … ) represent the kinetic terms for the forms not affected by the topology change. In the shrinking cycle limit $\tau_{2}\to 0^{+}$, the combination $H_{2}^{+}=H_{2}^{c}-\tau_{1}H_{2}^{c^{\prime}}$ becomes weakly couples, while $H_{2}^{-}=H_{2}^{c^{\prime}}$ becomes strongly coupled.
    As the 1-cycle $c$ shrinks along the bordism direction $\rho$ in the topology change process, the 1-forms associated with the angular coordinate along $c$, $\mathrm{d}\theta$, ceases to be harmonic, as the internal metric is warped with respect to $\rho$ (i.e., $\mathcal{B}_{3}$ is no longer locally $\Sigma_{g}$ times an interval). When decomposing $B_{2}=B_{1}\wedge\mathrm{d}\theta$, close to the topology change locus $\Delta\theta\neq 0$, in such a way that the 8d field $B_{1}$ is massed-up and the $U(1)$ 1-form symmetry is removed from the EFT. At the same time, the angular profile $\theta$ of the condensing tachyon (B.2), which is wound along it, acts as a Goldstone boson of the broken $U(1)$ symmetry, see Delgado:2023uqk . As for the gauge field associated to the dual $c^{\prime}$, fundamental strings charged under the $B_{2}$ field and winding around $c^{\prime}$ in opposite directions are seen as charge-anticharge pairs from the (7+1)-dimensional point of view, charged under $B_{1}^{+}$, where $H_{2}^{+}=\mathrm{d}B_{1}^{+}$. As $\tau_{2}\to 0^{+}$, from (4.2) the gauge coupling of these grows stronger, with these pairs experimenting a strong attractive force (think of tension full flux string in 8d), until they encounter and annihilate each other. This results in the winding charge along $c^{\prime}$ to be effectively 0 as $\tau_{2}\to 0$ and the $U(1)$ 1-form gauge symmetry is confined, thus not obstructing the handle decay. This is illustrated in Figure 7
    In the new EFT resulting on the other side of the topology change domain wall, the rank of the gauge group is two units lower, as the two $U(1)$ 1-forms associated to the two no-longer-present cycles have either been massed-up or classically confined Adams:2005rb , thus changing the gauge symmetry of the theory from $U(1)^{2g}$ to $U(1)^{2g-2}$.
@@
In any case, we find that after any of the $g$ or $g+1$ necessary topology changes in the bordism from Type IIB string theory on $\Sigma_{g}$ to nothing, the vacuum energy lowers and the rank of the gauge group becomes smaller, with the different changes being energetically favored.
@@
where we do not distinguish between the $\hat{\tau}_{2}\to\pm\infty$ limits which are equivalent for the potential, and as we will see now not necessarily the three scalars simultaneously blow up. From our analysis, there are the following (qualitatively) different topology changes:
  • $\Sigma_{g}\to\Sigma_{g-1}$: Here one of the 1-cycles shrinks to a point, i.e., $|\hat{\tau}_{2}|\to\infty$, while the overall volume is remains approximately constant. On the other hand, as we approach the D7-brane, the dilaton goes to a weak coupling limit, $\hat{\Phi}\to-\infty$, see section 18.5 from Blumenhagen:2013fgp . This results in $V(\varphi)\sim\exp\left(\sqrt{2}\hat{\Phi}+\sqrt{2}|\hat{\tau}_{2}|\right)\;$, with critical exponent $\delta=1$.
@@ 4.3 An example of non-minimal bordism
In the previous two examples studied, the bordisms considered saturated the Morse-Bott (3.13) and $\mathcal{B}_{n+1}$ homology (3.8) inequalities, in such a way that the number of topology changes was minimized to only those unavoidable. There are, however, some results in the literature in which smooth bordisms which are non-minimal, i.e., those inequalities are fulfilled but not saturated. One of such examples was commented upon in Figure 5(b) and section 3.3, Debray:2023yrs . For a more detailed instance, we will comment on the BoN construction described in GarciaEtxebarria:2020xsr , whose relevant ingredients we review now. The following $D$-dimensional (bosonic) effective action is considered, which for simplicity we present in string frame [^20] The Gauss-Bonnet invariant $\mathsf{R}_{\rm GB}^{2}$ transforms in a pretty involved way under the conformal rescaling required to move to Einstein frame, see eq. (III.4) from Dabrowski:2008kx . As we will not make use of the actual equations of motion, we simply present the effective action in order to present the field content.
@@
where $\Phi$ is the $D$-dimensional dilaton, $H_{3}=\mathrm{d}B_{2}$ and $\mathsf{R}_{\rm GB}^{2}=\mathcal{R}^{2}-4R_{MN}R^{MN}+R_{MNPQ}R^{MNPQ}$ is the Gauss-Bonnet invariant, which is topological for $D=4$ and gives second-order e.o.m. for higher $D$. For $\alpha>0$ both SUSY and the Dominant Energy Condition are broken. See GarciaEtxebarria:2020xsr for more details on the embedding of the above action in $D=6$ heterotic and the dual type I/IIB duals.
@@
The above effective action can be further compactified to $d=3$ on a 3-torus $\mathcal{C}_{3}=\mathbb{T}^{3}$ with SUSY preserving (i.e., periodic) boundary conditions. The bordism structure considered is $\mathsf{g}={\rm Spin}$, and since $\Omega_{3}^{\rm Spin}\simeq 0$ Anderson1967:SPIN ; McNamara:2019rup , there exists a Spin manifold $\mathcal{B}_{4}$ with $\mathcal{C}_{3}=\partial\mathcal{B}_{4}$, and thus from the discussion in section 2.1 BoN constructions should exist. Due to the Spin structure on $\mathcal{C}_{3}$, $\mathcal{B}_{4}$ cannot simply be $\mathbb{D}^{2}\times\mathbb{T}^{2}$, as we cannot just contract one of the 1-cycles to a point. Without the need of including stringy Spin-defects, a smooth solution is precisely given in Scorpan2005TheWW ; GarciaEtxebarria:2020xsr , with $\mathcal{B}_{4}$ given by elliptic fibration over $\mathbb{C}\simeq\mathbb{D}^{2}$ such that the $\mathbb{T}^{2}$-fiber pinches in 12 points over the disk. If these degeneration loci are isolated, then $\mathcal{B}_{4}$ is smooth. In Figure 8 a sketch of the $\mathcal{B}_{4}$ manifold is presented. By continuously deforming it such that the 12 pinching points occurs on the same point of the base $\mathbb{D}^{2}$, we can easy obtain $\vec{b}(\mathcal{B}_{4})=(1,1,12,0,0)$ (rather than $\vec{b}(\mathbb{D}^{2}\times\mathbb{T}^{2})=(1,2,1,0,0)$), with no torsion. Note that these Betti numbers fulfill but do not saturate the inequalities (3.8). In any case, Morse-Bott inequalities (3.13) are satisfied as
@@
where both the 12 pinching points have index 2 and the center of $\mathbb{D}^{2}$, where the 1-cycle closes over the $\mathbb{T}^{2}$ fiber, have index 2. Note that after each critical point, where a 1-cycle pinches and then grows again, the local topology of the solution is actually the same, $\mathbb{T}^{3}$, so that no actual topology change occurs. In any case, these types of critical points also enter in the Morse-Bott (3.13) and $\mathcal{B}_{n+1}$ homology (3.8) inequalities, with certainly the bordism into question being different than $\mathbb{T}^{2}\times\mathbb{D}^{2}$.
As explained in great detail in GarciaEtxebarria:2020xsr , each of the 12 degeneration points can be locally seen as a Taub-NUT/KK monopole solution Gross:1983hb ; Sorkin:1983ns , with the compact space smoothly pinching, so that could also consider this solution as a 12-centered BoN solution. Close to this degeneration points, there is a partial decompatification, with the elliptic fiber becoming large while the $\mathbb{S}^{1}$ basis shrinks (so that the overall volume remains finite). This results in an infinite distance limit in moduli space being proven. [^21] As explained in section 5.2.2 from GarciaEtxebarria:2020xsr , the associated KK tower is vital to cancel an a priori divergence in the metric around the degeneration points, such that the BoN solution remains smooth. From the EotW brane approach from section 2.2, there is no potential here for it to blow-up, so that $a=0$ and $V_{0}=0$, and the critical exponent is $\delta=2\sqrt{\frac{d-1}{d-2}}$. Again we refer to the extensive work of GarciaEtxebarria:2020xsr for the detailed expressions.
This example might feel a bit underwhelming, as it involves a topology for $\mathcal{B}_{4}$ much more complicated than $\mathbb{T}^{2}\times\mathbb{D}^{2}$, motivated by having a smooth bordism rather than using stringy defects such as those mentioned in footnote 6. Given the topology of $\mathcal{B}_{4}$, which as described in GarciaEtxebarria:2020xsr is determined by requiring 12 degeneration points of the elliptic fiber over the $\mathbb{D}^{2}$ base, one can use Morse-Bott inequalities to show that
[…]
where we denote $A\twoheadrightarrow B$ as a cycle $A$ shrinking over a critical locus $B$. For our example we have $121-b_{1}(\mathcal{B}_{4})+b_{2}(\mathcal{B}_{4})=$ 12, with only 12 1-cycles shrinking over points and a final shrinking $\mathbb{S}^{1}\twoheadrightarrow\mathbb{T}^{2}=\Sigma_{1}$, with no 2-cycles shrinking over a point. Note that, in this aspect, the number of critical points is the minimum needed, and other solutions to the above equations would result in a higher amount of them.
@@
[…]
@@
    $(b,c)=(1,0)$: Assume that after the intermediate $a$ topology changes in each of which two (dual) 2-cycles disappear we end up with a 4-manifold $\mathcal{X}_{4}$ with $\vec{b}(\mathcal{X}_{4})=(1,0,2\gamma,0,1)$ before the final topology change into nothing. By similar arguments as above, this final bordism $\mathcal{M}^{\prime}_{5}$ will have $\vec{b}(\mathcal{M}^{\prime}_{5})=(1,0,\gamma,\beta^{\prime},0,0)$ and by Morse-Bott inequalities $\gamma=1-2g-\beta^{\prime}$, with $\beta^{\prime}\in\{0,1\}$. This results in $g=0$ and two additional cases:
@@
      $\beta^{\prime}=0$: Then $\gamma=1$ and $\mathcal{P}_{t}(\mathcal{X}_{4})=1+2t^{2}+t^{4}=(1+t^{2})^{2}$, and as such can be seen as a fibration $\mathbb{S}^{2}\hookrightarrow\mathcal{X}_{4}\to\mathbb{S}^{2}$, such as one of the Hirzebruch surfaces $H_{n}$ Hirzebruch1951 . Note that these have different topological invariants for distinct values of $n$, so on general they would correspond with different topologies of $\mathcal{B}_{5}$. We also have that $a=10+2\hat{a}$, with the extra $2\hat{a}=\beta\geq 0$ topology changes corresponding to creation and disappearance of pairs of 2-cycles.
@@
      $\beta^{\prime}=1$: Then $\gamma=0$ and $\mathcal{P}_{t}(\mathcal{X}_{4})=1+t^{4}$, so $\mathcal{X}_{4}$ is an homology sphere. Again these are not unique SUCIU1987103 , and could give rise to different bordism topologies. There will be $a=11+2\hat{a}$, with additional $2\hat{a}=\beta-1\geq 0$ topology changes corresponding to additional creation and disappearance of pairs of 2-cycles.
@@
[…]
In section 7.2 of Debray:2023yrs a different bordism from K3 to nothing was argued for. F-theory on K3 (seen as a $\mathbb{T}^{2}\hookrightarrow{\rm K3}\to\mathbb{CP}^{2}$ fibration) is considered, which in Sen’s weak coupling limit is dual to perturbative type IIB string theory on a $\mathbb{T}^{2}/\mathbb{Z}_{2}$ orbifold. The four resulting singularities, which are specified by the monodromy of a linking circle, are argued to be equivalent to four $\mathbb{T}^{2}$ with the same monodromy. This motivates the substitution of $\mathbb{T}^{2}/\mathbb{Z}_{2}$ by these genus-4 surface, which now allows for a smooth bordism to nothing as in section 4.2, and over which the F-theory elliptic fiber does not degenerate. The problem comes in that the process $(\mathbb{T}^{2}\hookrightarrow{\rm K3}\to\mathbb{CP}^{2})\to(\mathbb{T}^{2}\hookrightarrow\mathcal{X}_{4}\to\Sigma_{4})$ is not necessarily smooth, with this $\mathcal{X}_{4}$, which contains non-trivial 1-cycles, not being required as an intermediate step of the analysis considered previously in this subsection. Since the construction of a completely smooth bordism to nothing for K3 compactifications (which for F-theory constructions need to keep the elliptic fibration until the final topology change into nothing) is not yet a solved matter, we hope this short digression serves to bring some light into the matter.

### v2

@@ 4.2 Type IIB string theory on compact Riemann surfaces
[…]
@@
  • We now turn to the other term in the (4.14) action, corresponding to the gauge kinetic term for the resulting 2-from fields in the 8d EFT, $\mathcal{A}_{AB}(\vec{\varphi})H_{2}^{A}\wedge\star H_{2}^{B}$, with $A=1,\,\dots,\,2g$. As explained in Adams et al. (2005), for a 1-cycle $c$ shrinking, with dual $c^{\prime}$ , both taken to be decoupled from the rest, we have that [^21] Note that there is a different sign and $\tau_{2}$ factor in the expression when compared to that in Adams et al. (2005), though the conclusions are the same.
    $\mathbb{A}_{AB}H_{2}^{A}\wedge\star H_{2}^{B}\supseteq\frac{\star 1}{\tau_{2}}|H_{2}^{c}-\tau H_{2}^{c^{\prime}}|^{2}=\frac{1}{\tau_{2}}|H_{2}^{-}|^{2}+\tau_{2}|H_{2}^{+}|^{2}\,$, (4.23)
    where $\tau=\tau_{1}+i\tau_{2}$ is the factorized complex structure associated to $c$ and $c^{\prime}$ and where we assume that the kinetic terms for the additional forms are not affected by the topology change. In the shrinking cycle limit $\tau_{2}\to 0^{+}$, the combination $H_{2}^{+}=H_{2}^{c}-\tau_{1}H_{2}^{c^{\prime}}$ becomes weakly coupled, while $H_{2}^{-}=H_{2}^{c^{\prime}}$ becomes strongly coupled.
    As the 1-cycle $c$ shrinks along the bordism direction $\rho$ in the topology change process, the 1-forms associated with the angular coordinate along $c$, $\mathrm{d}\theta$, ceases to be harmonic, as the internal metric is warped with respect to $\rho$ (i.e., $\mathcal{B}_{3}$ is no longer locally $\Sigma_{g}$ times an interval). When decomposing $B_{2}=B_{1}\wedge\mathrm{d}\theta$, close to the topology change locus $\Delta\theta\neq 0$, in such a way that the 8d field $B_{1}$ is massed-up and the $U(1)$ 1-form symmetry is removed from the EFT. At the same time, the angular profile $\theta$ of the condensing tachyon (B.2), which is wound along it, acts as a Goldstone boson of the broken $U(1)$ symmetry, see Delgado (2024). As for the gauge field associated to the dual $c^{\prime}$, fundamental strings charged under the $B_{2}$ field and winding around $c^{\prime}$ in opposite directions are seen as charge-anticharge pairs from the (7+1)-dimensional point of view, charged under $B_{1}^{+}$, where $H_{2}^{+}=\mathrm{d}B_{1}^{+}$. As $\tau_{2}\to 0^{+}$, from (4.23) the gauge coupling of these grows stronger, with these pairs experimenting a strong attractive force (think of tension full flux string in 8d), until they encounter and annihilate each other. This results in the winding charge along $c^{\prime}$ to be effectively 0 as $\tau_{2}\to 0$ and the $U(1)$ 1-form gauge symmetry is confined, thus not obstructing the handle decay. This is illustrated in Figure 7.
    In the new EFT resulting on the other side of the topology change domain wall, the rank of the gauge group is two units lower, as the two $U(1)$ 1-forms associated to the two no-longer-present cycles have either been massed-up or classically confined Adams et al. (2005), thus changing the gauge symmetry of the theory from $U(1)^{2g}$ to $U(1)^{2g-2}$.
@@
Summing up, we find that after any of the $g$ or $g+1$ necessary topology changes in the bordism from Type IIB string theory on $\Sigma_{g}$ to nothing, the vacuum energy lowers and the rank of the gauge group becomes smaller, with the different changes being energetically favored.
@@
where we do not distinguish between the $\hat{\tau}_{2}\to+\infty$ and -$\infty$ limits, which are equivalent for the potential, and as we will see now not necessarily the three scalars simultaneously blow up. From our analysis, there are the following (qualitatively) different topology changes:
  • $\Sigma_{g}\to\Sigma_{g-1}$: Here one of the 1-cycles shrinks to a point, i.e., $|\hat{\tau}_{2}|\to\infty$, while the overall volume is remains approximately constant. On the other hand, as we approach the D7-brane, the dilaton goes to a weak coupling limit, $\hat{\Phi}\to-\infty$, see section 18.5 from Blumenhagen et al. (2013). This results in $V(\varphi)\sim\exp\left(\sqrt{2}\hat{\Phi}+\sqrt{2}|\hat{\tau}_{2}|\right)\;$, with critical exponent $\delta=1$.
@@ 4.3 An example of non-minimal bordism
In the previous examples studied, the bordisms considered saturated the Morse-Bott (3.13) and $\mathcal{B}_{n+1}$ homology (3.8) inequalities, in such a way that the number of topology changes was minimized to only those unavoidable. There are, however, some results in the literature in which smooth bordisms which are non-minimal, i.e., those inequalities are fulfilled but not saturated. One of such examples was commented upon in Figure 5(b) and section 3.3, Debray et al. (2023). For a more detailed instance, we will comment on the BoN construction described in García Etxebarria et al. (2020), whose relevant ingredients we review now. The following $D$-dimensional (bosonic) effective action is considered, which for simplicity we present in string frame [^22] The Gauss-Bonnet invariant $\mathsf{R}_{\rm GB}^{2}$ transforms in a pretty involved way under the conformal rescaling required to move to Einstein frame, see eq. (III.4) from Dabrowski et al. (2009). As we will not make use of the actual equations of motion, we simply present the effective action in order to present the field content.
@@
where $\Phi$ is the $D$-dimensional dilaton, $H_{3}=\mathrm{d}B_{2}$ and $\mathsf{R}_{\rm GB}^{2}=\mathcal{R}^{2}-4R_{MN}R^{MN}+R_{MNPQ}R^{MNPQ}$ is the Gauss-Bonnet invariant, which is topological for $D=4$ and gives second-order e.o.m. for higher $D$. For $\alpha>0$ both SUSY and the Dominant Energy Condition are broken. See García Etxebarria et al. (2020) for more details on the embedding of the above action in $D=6$ heterotic and the dual type I/IIB descriptions.
@@
The above effective action can be further compactified to $d=3$ on a 3-torus $\mathcal{C}_{3}=\mathbb{T}^{3}$ with SUSY preserving (i.e., periodic) boundary conditions. The bordism structure considered is $\mathsf{g}={\rm Spin}$, and since $\Omega_{3}^{\rm Spin}\simeq 0$ Anderson et al. (1967); McNamara and Vafa (2019), there exists a Spin manifold $\mathcal{B}_{4}$ with $\mathcal{C}_{3}=\partial\mathcal{B}_{4}$, and thus from the discussion in section 2.1 BoN constructions should exist. Due to the Spin structure on $\mathcal{C}_{3}$, $\mathcal{B}_{4}$ cannot simply be $\mathbb{D}^{2}\times\mathbb{T}^{2}$, as we cannot just contract one of the 1-cycles to a point. Without the need of including stringy Spin-defects, a smooth solution is precisely given in Scorpan (2005); García Etxebarria et al. (2020), with $\mathcal{B}_{4}$ given by elliptic fibration over $\mathbb{C}\simeq\mathbb{D}^{2}$ such that the $\mathbb{T}^{2}$-fiber pinches in 12 points over the disk. If these degeneration loci are isolated, then $\mathcal{B}_{4}$ is smooth. In Figure 8 a sketch of the $\mathcal{B}_{4}$ manifold is presented. By continuously deforming it such that the 12 pinching points occurs on the same point of the base $\mathbb{D}^{2}$, we can easily obtain $\vec{b}(\mathcal{B}_{4})=(1,1,12,0,0)$ (rather than $\vec{b}(\mathbb{D}^{2}\times\mathbb{T}^{2})=(1,2,1,0,0)$), with no torsion. Note that these Betti numbers fulfill but do not saturate the inequalities (3.8). In any case, Morse-Bott inequalities (3.13) are satisfied as
@@
where both the 12 pinching points have index 2 and the center of $\mathbb{D}^{2}$, where the 1-cycle closes over the $\mathbb{T}^{2}$ fiber, has index 2. Note that after each critical point, where a 1-cycle pinches and then grows again, the local topology of the solution is actually the same, $\mathbb{T}^{3}$, so that no actual topology change occurs. In any case, these types of critical points also enter in the Morse-Bott (3.13) and $\mathcal{B}_{n+1}$ homology (3.8) inequalities, with certainly the bordism into question being different than $\mathbb{T}^{2}\times\mathbb{D}^{2}$.
As explained in great detail in García Etxebarria et al. (2020), each of the 12 degeneration points can be locally seen as a Taub-NUT/KK monopole solution Gross and Perry (1983); Sorkin (1983), with the compact space smoothly pinching, so that could also consider this solution as a 12-centered BoN solution. Close to this degeneration points, there is a partial decompatification, with the elliptic fiber becoming large while the $\mathbb{S}^{1}$ basis shrinks (so that the overall volume remains finite). This results in an infinite distance limit in moduli space being proven. [^23] As explained in section 5.2.2 from García Etxebarria et al. (2020), the associated KK tower is vital to cancel an a priori divergence in the metric around the degeneration points, such that the BoN solution remains smooth. From the EotW brane approach from section 2.2, there is no potential here for it to blow-up, so that $a=0$ and $V_{0}=0$, and the critical exponent is $\delta=2\sqrt{\frac{d-1}{d-2}}$. Again we refer to the extensive work of García Etxebarria et al. (2020) for the detailed expressions.
This example might feel a bit underwhelming, as it involves a topology for $\mathcal{B}_{4}$ much more complicated than $\mathbb{T}^{2}\times\mathbb{D}^{2}$, motivated by having a smooth bordism rather than using stringy defects such as those mentioned in footnote 6. Given the topology of $\mathcal{B}_{4}$, which as described in García Etxebarria et al. (2020) is determined by requiring 12 degeneration points of the elliptic fiber over the $\mathbb{D}^{2}$ base, one can use Morse-Bott inequalities to show that
$\{\text{1-cycle $\twoheadrightarrow\{\rm pt.\}$}\}-\#\{\text{2-cycle $\twoheadrightarrow\{\rm pt.\}$}\}+\left\{\begin{array}[]{ll}2(1-g)&\text{if $\mathbb{S}^{1}\twoheadrightarrow\Sigma_{g}$ end}\\ 0&\text{otherwise}\end{array}\right.=1-b_{1}(\mathcal{B}_{4})+b_{2}(\mathcal{B}_{4})\;,$
where we denote $A\twoheadrightarrow B$ as a cycle $A$ shrinking over a critical locus $B$. For our example we have $1-b_{1}(\mathcal{B}_{4})+b_{2}(\mathcal{B}_{4})=12$, with only 12 1-cycles shrinking over points and a final shrinking $\mathbb{S}^{1}\twoheadrightarrow\mathbb{T}^{2}=\Sigma_{1}$, with no 2-cycles shrinking over a point. Note that, in this aspect, the number of critical points is the minimum needed, and other solutions to the above equations would result in a higher amount of them.
@@
We end this section with a comment on the bordism structure for a more complicated internal manifold than the Riemann surfaces from section 4.2: the 4-dimensional K3 surface. Considering the $\mathsf{g}={\rm Spin}$ structure, we have that $\Omega_{4}^{\rm Spin}\simeq\mathbb{Z}\simeq\langle{\rm K3}\rangle$ Anderson et al. (1967), this is, precisely generated by K3, which carries negative unit charge under the conserved current $J_{p_{1}}=\frac{1}{48}p_{1}(R)=\frac{1}{48\cdot 8\pi^{2}}{\rm tr}(R\wedge R)$, with $p_{1}$ the first Pontriaguin class of the curvature 2-form $R$. How this $U(1)$ symmetry is killed in different string theories is discussed in detail in McNamara and Vafa (2019), to which we refer for more details. For our purposes, type IIA/M-theory can be compactified on pin+-manifolds Witten (2016), on which the bordism is broken to $\Omega_{4}^{\rm Spin}\simeq\mathbb{Z}\to\Omega^{\rm pin+}_{4}\simeq\mathbb{Z}_{16}\simeq\langle\mathbb{RP}^{4}\rangle$, which is then broken by MO5-planes Dasgupta and Mukhi (1996); Witten (1996). As for type IIB string theory, this is partly broken by non-trivial elliptic fibrations in F-theory, while, as argued in McNamara and Vafa (2019), there must be some new, non-supersymmetric defect with spatial dimensionality 6 that kills the remaining the remaining class. In order to obtain smooth bordisms, we will assume that defects are located far from the tip of the bordism, in such a way that our K3 surface can be taken belong to the trivial class of the spin bordism. As in section 4.2, we will only care about flux potentials and the location of the branes needed to eat-up the associated charges.
Before this, it is necessary to obtain information about the topology of the bordism $\mathcal{B}_{5}$ with ${\rm K3}=\partial\mathcal{B}_{5}$. As K3 has Poincaré polynomial $\mathsf{P}_{t}({\rm K3})=1+22t^{2}+t^{4}$ (i.e. we only have 2-cycles and the whole 4-manifold to shrink). Using (3.8) and (3.9) we find that any bordism $\mathcal{B}_{5}$ for it must have Betti numbers $\vec{b}(\mathcal{B}_{5})=(1,\alpha,11,\beta,0,0)$, with $\alpha,\,\beta\in\mathbb{Z}_{\geq 0}$. Without creating new cycles, we have that the following possible topology changes to (A) 2-cycles shrinking to a point, (B) a 2-cycle shrinking over a compact 2-cycle of genus $g$, and (C) a 4-cycle shrinking to a point. The last two possibilities make 4 internal dimensions disappear, and as such correspond to the final topology change. Respectively denoting the number of each type by $a\in\mathbb{Z}_{\geq 0}$ and $b,\;c\in\{0,1\}$, we will then have that the LHS in (3.12) reads as
@@
    $(b,c)=(1,0)$: Assume that after the intermediate $a$ topology changes ,in each of which two (dual) 2-cycles disappear, we end up with a 4-manifold $\mathcal{X}_{4}$ with $\vec{b}(\mathcal{X}_{4})=(1,0,2\gamma,0,1)$ before the final topology change into nothing. By similar arguments as above, this final bordism $\mathcal{M}^{\prime}_{5}$ will have $\vec{b}(\mathcal{M}^{\prime}_{5})=(1,0,\gamma,\beta^{\prime},0,0)$ and by Morse-Bott inequalities $\gamma=1-2g-\beta^{\prime}$, with $\beta^{\prime}\in\{0,1\}$. This results in $g=0$ and two additional cases:
@@
      $\beta^{\prime}=0$: Then $\gamma=1$ and $\mathcal{P}_{t}(\mathcal{X}_{4})=1+2t^{2}+t^{4}=(1+t^{2})^{2}$, and as such can be seen as a fibration $\mathbb{S}^{2}\hookrightarrow\mathcal{X}_{4}\to\mathbb{S}^{2}$, such as one of the Hirzebruch surfaces $H_{n}$ Hirzebruch (1951). Note that these have different topological invariants for distinct values of $n$, so on general they would correspond to different topologies of $\mathcal{B}_{5}$. We also have that $a=10+2\hat{a}$, with the extra $2\hat{a}=\beta\geq 0$ topology changes corresponding to creation and disappearance of pairs of 2-cycles.
@@
      $\beta^{\prime}=1$: Then $\gamma=0$ and $\mathcal{P}_{t}(\mathcal{X}_{4})=1+t^{4}$, so $\mathcal{X}_{4}$ is an homology sphere. Again these are not unique Suciu (1987), and could give rise to different bordism topologies. There will be $a=11+2\hat{a}$, with additional $2\hat{a}=\beta-1\geq 0$ topology changes corresponding to additional creation and disappearance of pairs of 2-cycles.
@@
One could perform a similar analysis to that of section 4.2 to compactifications on K3. In this case, as $b_{k}({\rm K3})=0$ for odd $k$, only even fluxes contribute to the flux potential, so type IIA compactifications will be more interesting than type IIB, with D4- and wrapped D6-branes (with bound D2-branes from the dual 2-cycle) serving as domain walls between topology changes in the presence of flux along the whole K3 or the different 2-cycles. The analysis of the lower-dimensional gauge fields fate after topology change is analogous to those in section 4.2, so we will not comment on it further.
In section 7.2 of Debray et al. (2023) a different bordism from K3 to nothing was argued for. F-theory on K3 (seen as a $\mathbb{T}^{2}\hookrightarrow{\rm K3}\to\mathbb{CP}^{2}$ fibration) is considered, which in Sen’s weak coupling limit is dual to perturbative type IIB string theory on a $\mathbb{T}^{2}/\mathbb{Z}_{2}$ orbifold. The four resulting singularities, which are specified by the monodromy of a linking circle, are argued to be equivalent to four $\mathbb{T}^{2}$ with the same monodromy. This motivates the substitution of $\mathbb{T}^{2}/\mathbb{Z}_{2}$ by these genus-4 surface, which now allows for a smooth bordism to nothing as in section 4.2, and over which the F-theory elliptic fiber does not degenerate. The problem comes in that the process $(\mathbb{T}^{2}\hookrightarrow{\rm K3}\to\mathbb{CP}^{2})\to(\mathbb{T}^{2}\hookrightarrow\mathcal{X}_{4}\to\Sigma_{4})$ is not necessarily smooth, with this $\mathcal{X}_{4}$, which contains non-trivial 1-cycles, not being required as an intermediate step of the analysis considered previously in this subsection. Since the construction of a completely smooth bordism to nothing for K3 compactifications (which for F-theory constructions need to keep the elliptic fibration until the final topology change into nothing) is not yet a solved matter, we hope this short digression serves to bring some light into the matter.